# Hierarchical Context Alignment with Disentangled Geometric and Temporal Modeling for Semantic Occupancy Prediction

### v1

@@ 3.7 Training Objectives
where several $\lambda$s are balancing coefficients.
TABLE I: Quantitative comparison with the state-of-the-art camera-based semantic scene completion methods on the SemanticKITTI validation set. The “S-T”, “S” and “M” denote temporal stereo images, single-frame stereo images, and single-frame monocular images, respectively. The top two performers are marked bold and underline.
@@
TABLE II: Quantitative results with the state-of-the-art semantic scene completion methods on the SemanticKITTI test set. The “S-T”, “S” and “M” denote temporal stereo images, single-frame stereo images, and single-frame monocular images, respectively. The top two performers are marked bold and underline.
@@
TABLE III: Quantitative comparison with the state-of-the-art semantic scene completion methods on the NuScenes-Occupancy validation set. The top two performers are marked bold and underline. The “L”, “M”, “M-D” and “M-T” denote LiDAR inputs, monocular images, monocular images with depth maps and temporal monocular images, respectively. The LiDAR points are projected and densified to generate the depth maps.
Methods Hi-SOP(M) HTCL-M JS3C-Net LMSCNet 3DSketch AICNet [3] TPVFormer MonoScene
@@
4 Experiment
4.1 Datasets and Metrics
@@
SemanticKITTI. The SemanticKITTI dataset [25] includes 22 outdoor scenes characterized by LiDAR scans and stereo images. The ground truth is structured into 256$\times$256$\times$32 voxel grids, with each voxel measuring 0.2m in all dimensions and annotated with 21 semantic classes (19 semantic, 1 free, and 1 unknown). Consistent with prior studies [16, 5], we divided the dataset into 10 training scenes, 1 validation scene, and 1 test scene. We evaluated our method using both stereo (HTCL-S) and monocular images Hi-SOP(M) on the SemanticKITTI.
@@
NuScenes. The NuScenes dataset [27] is an autonomous driving dataset collected in Boston and Singapore. It comprises 1,000 driving sequences across various environments, with each sequence lasting approximately 20 seconds. Keyframes are annotated at a rate of 2Hz with 3D bounding boxes. The Panoptic NuScenes dataset [84] extends the original NuScenes dataset by providing annotations for LiDAR semantic segmentation. Following previous works [6, 17], we utilize sparse LiDAR point supervision for 3D semantic occupancy prediction. We divided the dataset into training, validation, and testing splits containing 700, 150, and 150 scenes, respectively. Note that our monocular-based approach of Hi-SOP(M) is exclusively applied on the NuScenes dataset due to the absence of stereo images.
@@
NuScenes-Occupancy. The NuScenes-Occupancy dataset [26] is an extension of the NuScenes dataset [27], which provides dense semantic occupancy annotations for 850 scenes comprising 34,000 keyframes with 360-degree LiDAR scans. We divide the dataset into 28,130 training frames and 6,019 validation frames as described in [26]. Each frame includes 400K occupied voxels labeled with 17 semantic classes. We exclusively apply our monocular-based Hi-SOP(M) on the OpenOccupancy dataset as on the NuScenes dataset.
@@
Evaluation Metrics. Following previous works [16, 5], we adopt the mean Intersection over Union (mIoU) as the primary metric for evaluating the Semantic Scene Completion (SSC) task and the LiDAR semantic segmentation task. Additionally, the Intersection over Union (IoU) metric is used to evaluate the performance of the class-agnostic scene completion (SC) task. For the evaluation of LiDAR semantic segmentation results, the LiDAR points are only used to query their corresponding semantic logits from the predicted 3D semantic occupancy volume following [6, 17].
4.2 Experimental Setup
@@
Following standard practices [16, 17, 5], we initialize the UNet backbone with pre-trained weights from EfficientNetB7 [64]. By default, the model takes the current and previous three image frames as inputs. We implement our model on PyTorch with a batch size of 4 and train the model for 24 epochs using the AdamW optimizer [85]. The learning rate is set at $1\times 10^{-4}$, with a weight decay of 0.01.
4.3 Main Results
4.3.1 Quantitative Comparison
@@
As detailed in Table I and Table II, we conducted a comparison analysis of our proposed method against existing best methods on the SemanticKITTI dataset, including VoxFormer  [5], OccFormer [17], SurroundOcc [7], TPVFormer [6], and MonoScene [16]. VoxFormer-T is a temporal baseline utilizing the current and previous four frames as inputs. Our method demonstrates superior performance, surpassing VoxFormer-T by 4.84 mIoU on the SemanticKITTI validation set and 4.08 mIoU on the SemanticKITTI test set, with fewer historical inputs (3 vs. 4).
Further quantitative evaluations on the NuScenes-Occupancy validation set are presented in Table III. For depth map generation required by AICNet [3] and 3DSketch [86], LiDAR points are projected and densified following [26]. Despite the inherent advantage of LiDAR in IoU measurements due to its accurate 3D geometrical data, our method outperforms all competing methods in terms of mIoU, including those based on LiDAR like LMSCNet [14] and JS3C-Net [34]. This demonstrates the robustness and effectiveness of Hi-SOP in the semantic occupancy prediction task.
The quantitative results on the NuScenes validation set are presented in Table IV. We compare our method with state-of-the-art camera-based methods of OccFormer [17] and TPVFormer [6], and LiDAR-based methods of SalsaNext [87], PolarNet [88] and RangeNet++ [89]. Despite LiDAR’s inherent advantage in accurate 3D geometric measurements, our method outperforms all the other methods in terms of mIoU.
@@ 4.3.2 Qualitative Comparison
Figure 9 presents the qualitative comparison between our proposed method and VoxFormer [5] on the SemanticKITTI validation set. As we can see from the figure, the real-world scenes are inherently complex, and the sparsity of the annotated ground truth presents significant challenges in fully reconstructing semantic scenes from limited visual cues. Our method surpasses VoxFormer in capturing a more complete and accurate layout of the scenery, as illustrated by the crossroads in the first and third rows. Additionally, our method effectively infers the scenery beyond the camera’s field of view, notably in shadowed areas shown in the first and fifth rows, and exhibits marked improvements in handling dynamic objects, such as trucks in the second and sixth rows.
@@
The qualitative results on the NuScenes validation set are presented in Figure 11. Following previous works [6], we use only RGB images as input, while the LiDAR points are only used to query their features and for supervision in the training phase. Our proposed method generates more accurate semantic labels compared with the results from TPVFormer [6].
4.3.3 Temporal Stereo Variants Evaluation.
To ensure a fair and comprehensive comparison, we have implemented temporal stereo variants of baseline models as detailed in Table V. Following the approach of VoxFormer-T, we utilize stacked temporal stereo images as inputs, creating variants of MonoScene [16], TPVFormer [6], and OccFormer [17]. It is important to note that whereas VoxFormer-T [5] originally utilizes four previous frames, the stereo variants we developed employ only three previous frames, aligning with our method. As demonstrated in the table, our approach consistently achieves superior performance using the same temporal inputs.
TABLE VI: Ablation study for different architectural components on the SemanticKITTI validation set. The full names of different components are in Sec 4.4.
@@ 4.4 Ablation Study
Effect of GCL. The ablation results for Geometric Confidence-aware Lifting (GCL) are presented in the second row of Table VI. As shown in the table, replacing the typical lifting process as previous work [30, 17] with our proposed Geometric Confidence-aware Lifting increases the IoU and mIoU by 2.39 and 1.25, respectively. We attribute such improvements to the explicit geometric modeling with depth distribution confidence.
@@
Module Quantity Setting. We conduct ablation studies of quantity setting for the Multi-group Context Generation and the Multi-level Deformable Block, as presented in the ninth row of Table VII. As introduced in Section 3.4, we employ multiple groups of contextual features to facilitate diverse independent similarity learning. The results in Table VII demonstrate that leveraging 3 contextual groups yields a significant performance improvement, while employing more groups (5 groups) leads to a relatively slight improvement. Similarly, the enhancement of utilizing more feature levels (5 levels) in the Multi-level Deformable Block is also relatively minor. Therefore, considering the time consumption and parameter efficiency, we adopt 3 contextual groups in the Multi-group Context Generation and 3 feature levels in the Multi-level Deformable Block as the default settings.
@@
Temporal Inputs. We evaluated the performance of semantic occupancy prediction and the associated computational times using varying numbers of temporal inputs, as outlined in Table VIII. The results indicate that the marginal gains in effectiveness when using more than three previous frames are minimal compared to the increase in computational time. Therefore, we have chosen three frames as our standard configuration to achieve an optimal balance between efficiency and effectiveness.
5 Conclusion

### v2

@@ III-G Training Objectives
[…]
TABLE I: Quantitative comparison with the state-of-the-art camera-based semantic occupancy prediction methods on the SemanticKITTI validation set. The “S-T”, “S” and “M” denote temporal stereo images, single-frame stereo images, and single-frame monocular images, respectively. The top two performers are marked bold and underline.
@@
TABLE II: Quantitative results with the state-of-the-art semantic occupancy prediction methods on the SemanticKITTI test set. The “S-T”, “S” and “M” denote temporal stereo images, single-frame stereo images, and single-frame monocular images, respectively. The top two performers are marked bold and underline.
@@
TABLE III: Quantitative comparison with the state-of-the-art semantic occupancy prediction methods on the NuScenes-Occupancy validation set. The top two performers are marked bold and underline. The “L”, “M”, “M-D” and “M-T” denote LiDAR inputs, monocular images, monocular images with depth maps and temporal monocular images, respectively. The LiDAR points are projected and densified to generate the depth maps.
Methods Hi-SOP(M) HTCL-M JS3C-Net LMSCNet 3DSketch AICNet [li2021anisotropic] TPVFormer MonoScene
@@
IV Experiment
IV-A Datasets and Metrics
@@
SemanticKITTI. The SemanticKITTI dataset [behley2019semantickitti] includes 22 outdoor scenes characterized by LiDAR scans and stereo images. The ground truth is structured into $\times256$$\times$32 voxel grids, with each voxel measuring 0.2m in all dimensions and annotated with 21 semantic classes (19 semantic, 1 free, and 1 unknown). Consistent with prior studies [cao2022monoscene, li2023voxformer], we divided the dataset into 10 training scenes, 1 validation scene, and 1 test scene. We evaluated our method using both stereo (HTCL-S) and monocular images Hi-SOP(M) on the SemanticKITTI.
@@
NuScenes. The NuScenes dataset [caesar2020nuscenes] is an autonomous driving dataset collected in Boston and Singapore. It comprises 1,000 driving sequences across various environments, with each sequence lasting approximately 20 seconds. Keyframes are annotated at a rate of 2Hz with 3D bounding boxes. The Panoptic NuScenes dataset [fong2022panoptic] extends the original NuScenes dataset by providing annotations for LiDAR semantic segmentation. Following previous works [huang2023tri, zhang2023occformer], we utilize sparse LiDAR point supervision for 3D semantic occupancy prediction. We divided the dataset into training, validation, and testing splits containing 700, 150, and 150 scenes, respectively. Note that our monocular-based approach of Hi-SOP(M) is exclusively applied on the NuScenes dataset due to the absence of stereo images.
@@
NuScenes-Occupancy. The NuScenes-Occupancy dataset [wang2023openoccupancy] is an extension of the NuScenes dataset [caesar2020nuscenes], which provides dense semantic occupancy annotations for 850 scenes comprising 34,000 keyframes with 360-degree LiDAR scans. We divide the dataset into 28,130 training frames and 6,019 validation frames as described in [wang2023openoccupancy]. Each frame includes 400K occupied voxels labeled with 17 semantic classes. We exclusively apply our monocular-based Hi-SOP(M) on the OpenOccupancy dataset as on the NuScenes dataset.
@@
Evaluation Metrics. Following previous works [cao2022monoscene, li2023voxformer], we adopt the mean Intersection over Union (mIoU) as the primary metric for evaluating the Semantic Scene Completion (SSC) task and the LiDAR semantic segmentation task. Additionally, the Intersection over Union (IoU) metric is used to evaluate the performance of the class-agnostic scene completion (SC) task. For the evaluation of LiDAR semantic segmentation results, the LiDAR points are only used to query their corresponding semantic logits from the predicted 3D semantic occupancy volume following [huang2023tri, zhang2023occformer].
IV-B Experimental Setup
@@
Following standard practices [cao2022monoscene, zhang2023occformer, li2023voxformer], we initialize the UNet backbone with pre-trained weights from EfficientNetB7 [tan2019efficientnet]. By default, the model takes the current and previous three image frames as inputs. We implement our model on PyTorch with a batch size of 4 and train the model for 24 epochs using the AdamW optimizer [loshchilov2017decoupled]. The learning rate is set at $1\times 10^{-4}$, with a weight decay of 0.01.
IV-C Main Results
IV-C1 Quantitative Comparison
@@
As detailed in Table I and Table II, we conducted a comparison analysis of our proposed method against existing best methods on the SemanticKITTI dataset, including VoxFormer [li2023voxformer], OccFormer [zhang2023occformer], SurroundOcc [wei2023surroundocc], TPVFormer [huang2023tri], and MonoScene [cao2022monoscene]. VoxFormer-T is a temporal baseline utilizing the current and previous four frames as inputs. Our method demonstrates superior performance, surpassing VoxFormer-T by 4.84 mIoU on the SemanticKITTI validation set and 4.08 mIoU on the SemanticKITTI test set, with fewer historical inputs (3 vs. 4).
Further quantitative evaluations on the NuScenes-Occupancy validation set are presented in Table III. For depth map generation required by AICNet [li2021anisotropic] and 3DSketch [3d-sketch], LiDAR points are projected and densified following [wang2023openoccupancy]. Despite the inherent advantage of LiDAR in IoU measurements due to its accurate 3D geometrical data, our method outperforms all competing methods in terms of mIoU, including those based on LiDAR like LMSCNet [roldao2020lmscnet] and JS3C-Net [yan2021sparse]. This demonstrates the robustness and effectiveness of Hi-SOP in the semantic occupancy prediction task.
The quantitative results on the NuScenes validation set are presented in Table IV. We compare our method with state-of-the-art camera-based methods of OccFormer [zhang2023occformer] and TPVFormer [huang2023tri], and LiDAR-based methods of SalsaNext [cortinhal2020salsanext], PolarNet [zhang2020polarnet] and RangeNet++ [milioto2019rangenet++]. Despite LiDAR’s inherent advantage in accurate 3D geometric measurements, our method outperforms all the other methods in terms of mIoU.
@@ IV-C2 Qualitative Comparison
Figure 9 presents the qualitative comparison between our proposed method and VoxFormer [li2023voxformer] on the SemanticKITTI validation set. As we can see from the figure, the real-world scenes are inherently complex, and the sparsity of the annotated ground truth presents significant challenges in fully reconstructing semantic scenes from limited visual cues. Our method surpasses VoxFormer in capturing a more complete and accurate layout of the scenery, as illustrated by the crossroads in the first and third rows. Additionally, our method effectively infers the scenery beyond the camera’s field of view, notably in shadowed areas shown in the first and fifth rows, and exhibits marked improvements in handling dynamic objects, such as trucks in the second and sixth rows.
@@
The qualitative results on the NuScenes validation set are presented in Figure 11. Following previous works [huang2023tri], we use only RGB images as input, while the LiDAR points are only used to query their features and for supervision in the training phase. Our proposed method generates more accurate semantic labels compared with the results from TPVFormer [huang2023tri].
IV-C3 Temporal Stereo Variants Evaluation.
To ensure a fair and comprehensive comparison, we have implemented temporal stereo variants of baseline models as detailed in Table V. Following the approach of VoxFormer-T, we utilize stacked temporal stereo images as inputs, creating variants of MonoScene [cao2022monoscene], TPVFormer [huang2023tri], and OccFormer [zhang2023occformer]. It is important to note that whereas VoxFormer-T [li2023voxformer] originally utilizes four previous frames, the stereo variants we developed employ only three previous frames, aligning with our method. As demonstrated in the table, our approach consistently achieves superior performance using the same temporal inputs.
TABLE VI: Ablation study for different architectural components on the SemanticKITTI validation set. The full names of different components are in Sec IV-D.
@@ IV-D Ablation Study
Effect of GCL. The ablation results for Geometric Confidence-aware Lifting (GCL) are presented in the second row of Table VI. As shown in the table, replacing the typical lifting process as in previous work [philion2020lift, zhang2023occformer] with our proposed Geometric Confidence-aware Lifting increases the IoU and mIoU by 2.39 and 1.25, respectively. We attribute such improvements to the explicit geometric modeling with depth distribution confidence.
@@
Module Quantity Setting. We conduct ablation studies of quantity setting for the Multi-group Context Generation and the Multi-level Deformable Block, as presented in the ninth row of Table VII. As introduced in Section III-D, we employ multiple groups of contextual features to facilitate diverse independent similarity learning. The results in Table VII demonstrate that leveraging 3 contextual groups yields a significant performance improvement, while employing more groups (5 groups) leads to a relatively slight improvement. Similarly, the enhancement of utilizing more feature levels (5 levels) in the Multi-level Deformable Block is also relatively minor. Therefore, considering the time consumption and parameter efficiency, we adopt 3 contextual groups in the Multi-group Context Generation and 3 feature levels in the Multi-level Deformable Block as the default settings.
@@
Temporal Inputs. The effect of using a different number of temporal frames in terms of prediction performance and inference speed are outlined in Table VIII. The results in indicate that the marginal gains in effectiveness when using more than three previous frames are minimal compared to the increase in computational time. Therefore, we have chosen three frames as our standard configuration to achieve an optimal balance between efficiency and effectiveness.
V Conclusion
@@ Acknowledgments
[Uncaptioned image] Bohan Li received the B.E. degree from the School of Control Engineering, Northeastern University (NEU), Shenyang, China, in 2019. He received the M.E. degree from the School of Control Science and Engineering, South China University of Technology (SCUT), Guangzhou, China, in 2022. He is currently pursuing the Ph.D. degree in Shanghai Jiao Tong University (SJTU) and Eastern Institute of Technology (EIT). His research interests include 3D visual perception, robotics, and multi-modality content generation.
[Uncaptioned image] Jiajun Deng is a research fellow at the University of Adelaide, Australian Institute for Machine Learning. He received his Ph.D. degree (2021) and a B.E degree (2016) from the department of Electrical Engineering and Information Science at the University of Science and Technology of China. He served as a guest editor of the Special Issue “Pre-trained Models for Multi-modality Understanding” of IEEE Transactions on Multimedia, in 2023. He also served as the Area Chair for ACM Multimedia, in 2024. His research interests include computer vision, multi-modality understanding, and embodied AI.
[Uncaptioned image] Yasheng Sun received the B.E. degree from Nanjing University of Aeronautics and Astronautics, Nanjing, China, in 2017. He received the M.E. degree from the School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai, China, in 2020. He received the Ph.D. degree in Computer Science from the School of Computing, Tokyo Institute of Technology, Japan, in 2024. His current research interest includes cross-modal generation, 3D generative model, stable diffusion model and its application in computer vision.
[Uncaptioned image] Xiaofeng Wang received the B.E. degree from the School of Automation, Nanjing University of Science and Technology (NJUST), Nanjing, China, in 2020. He is currently pursuing the Ph.D. degree in Institute of Automation, Chinese Academy of Science (CASIA), Beijing, China. His current research areas include 3D perception and video generation. He has co-authored 10+ journal and conference papers mainly on computer vision autonomous-driving problems, including CVPR, ECCV, ICCV, AAAI, and ICLR.
[Uncaptioned image] Xin Jin has been a tenure track Assistant Professor with the Eastern Institute of Technology (EIT), Ningbo, China. He is also a Researcher at the Ningbo Institute of Digital Twin. He received his Ph.D. degree in Electronic Engineering and Information Science from the University of Science and Technology of China (USTC). His research interests include computer vision, intelligent media computing, and deep learning. He has over 10 granted patent applications, around 40 publications, and over 3,500 Google citations. He is an IEEE member, and reviewer of IEEE Transactions on Image Processing (TIP), IEEE Transactions on Multimedia (TMM), and IEEE Transactions on Circuits and Systems for Video Technology (TCSVT).
[Uncaptioned image] Wenjun Zeng (Fellow, IEEE) received the B.E. degree from Tsinghua University, Beijing, China, in 1990, the M.S. degree from the University of Notre Dame, Notre Dame, IN, USA, in 1993, and the Ph.D. degree from Princeton University, Princeton, NJ, USA, in 1997. He has been a Chair Professor and the Vice President for Research at the Eastern Institute for Advanced Study (EIAS) / Eastern Institute of Technology (EIT), Ningbo, China, since October 2021. He is also the founding Executive Director of the Ningbo Institute of Digital Twin. He was a Sr. Principal Research Manager and a member of the Senior Leadership Team at Microsoft Research Asia, Beijing, from 2014 to 2021, where he led the video analytics research empowering the Microsoft Cognitive Services, Azure Media Analytics Services, Office, and Windows Machine Learning. He was with University of Missouri, Columbia, MO, USA from 2003 to 2016, most recently as a Full Professor. Prior to that, he had worked for PacketVideo Corp., Sharp Labs of America, Bell Labs, and Panasonic Technology. He has contributed significantly to the development of international standards (ISO MPEG, JPEG2000, and OMA). Dr. Zeng is on the Editorial Board of the International Journal of Computer Vision. He was an Associate Editor-in-Chief of the IEEE Multimedia Magazine and an Associate Editor of the IEEE TRANSACTIONS ON CIRCUITS AND SYSTEMS FOR VIDEO TECHNOLOGY, IEEE TRANSACTIONS ON INFORMATION FORENSICS AND SECURITY, and IEEE TRANSACTIONS ON MULTIMEDIA (TMM). He was on the Steering Committee of IEEE TRANSACTIONS ON MOBILE COMPUTING and IEEE TMM. He served as the Steering Committee Chair of IEEE ICME in 2010 and 2011, and has served as the General Chair or TPC Chair for several IEEE conferences (e.g., ICME’2018, ICIP’2017). He was the recipient of several best paper awards.